Compute the value of the discriminant and give the number of real solutions to the quadratic equation:
3x^2 - 7x + 9 = 0

Answer by fractalier(2101)

The discriminant for
3x^2 - 7x + 9 = 0
is found by b^2 - 4ac
Here that is
(-7)^2 - 4(3)(9) =
49 - 108 =
-59
Roots are complex, and no real roots exist.